OpenStudy (anonymous):

Evaluate the function H(x) = 18 – x, when x = –6. A.108 B.12 C.–24 D.24

Mehek (mehek14):

H(x) = 18 - (-6) = 18 + 6 = ?

OpenStudy (anonymous):

b or d ? I know it's one of those

Mehek (mehek14):

wut's 18 + 6?

OpenStudy (anonymous):

24

Mehek (mehek14):

that's the answer

OpenStudy (anonymous):

Thank you

Mehek (mehek14):

yw :)